# Discovery of Probabilistic Dirichlet-to-Neumann Maps on Graphs

### v1

@@ 5.2 Subsurface fracture network problem
Discrete fracture networks serve as preferential pathways for fluid flow and heat transport in subsurface systems. Understanding the dynamics of such systems is essential for applications such as carbon sequestration, geothermal energy, and hydraulic fracturing. However, observations of such systems are typically both sparse and noisy, rendering traditional approaches insufficient for proper study. Surrogate modeling with deep neural networks has been proposed as one potential approach to studying fracture network properties [35]. In the current work, we explore Dirichlet-to-Neumann maps on graphs as an alternative approach.
We use the synthetic 2-D fracture network dataset of [35], generated by a physics-based Darcy-flow solver. The fracture geometry is generated according to the procedure detailed in [9, 35], and is characterized by the percolation parameter, $p=9$, and aperture, $a=1.5$. Fracture segments are designated as edges, and fracture intersections and terminations are designated as vertices. The resulting graph, depicted in Figure 3, has $V=107$ vertices and $E=130$ edges, of which 17 are designated boundary boundary entities (nine injection sites with high hydraulic head and eight extraction sites with low hydraulic head).
@@ 5.3 Arterial blood flow problem
The high computational cost of building high fidelity models with full field patient-specific data makes arterial blood flow another appealing test case for this approach. Arterial blood flow is a challenging test case because it exhibits both nonlinear dynamics and potentially complicated geometry. Our dataset consists of arterial graphs generated as training data for the GNN reduced order model developed in [30]. Data generation involves first running the 3D hemodynamic simulation SimVascular [38], then converting the data to a 1D representation and ultimately into a graph. The full data generation process is described in Section 4 of [30].
@@
Because the approach described in [30] calculated the artery aperture using a cutting-plane method, the dataset can be nonphysical (and hence, non-conservative) in areas near intersections, since a single plane may cut more than one arterial branch. For this reason, when we consider results on the interior of the graph, we focus on the “branch” vertices where this data fidelity issue does not arise (highlighted in Figure 5 Panel B).
@@
The cutting-plane artifact in [30] breaks local mass conservation near junctions; our model, by construction, enforces a globally conservative flow field that minimizes the RKHS norm subject to the boundary data. Consequently, the largest discrepancies between the predictions and truth appear exactly in those non-conservative regions, as visualized by the blue and red shading in the rightmost panel of Figure 6. For branch vertices—where the true data are conservative—errors are small and are captured by the estimated uncertainty bound. This is because our approach constructs a conservative model obeying the boundary data regardless of whether the ground truth data is actually locally conservative everywhere. It is also important to note that we expect our estimated uncertainty bound to underestimate the error on the interior due to the fact that the bound assumes the GP input values on the interior are accurate, when they are actually inferred estimates (see Section 4.4).
@@ 6.1 Future work
We assumed small Gaussian noise in our observations. The effect of this assumption is the regularization of the kernel matrix with a small diagonal term $\sigma_{\epsilon}^{2}I$, referred to as a “nugget.” For simplicity, we have assumed a uniform value for $\sigma_{\epsilon}^{2},$ but in reality the spectral properties of the different edges of our graph (and thus the different blocks of the kernel matrix) may differ widely. An adaptive nugget term, such as that proposed by [6], could more optimally balance accuracy and stability across the different edges of the graph.
Our current treatment does not account for the uncertainty stemming from the inferred input values to the GPs. We assume that the interior vertex values are unknown and estimated via the procedure described in Section 4.4, but the error estimates provided in Section 4.5 do not account for this. To incorporate this additional source of uncertainty, future work could explore linearizing the problem around the MAP, for example via a Laplace approximation as in [7].
Finally, future work could explore learning the kernel $K(\cdot,\cdot)$ that minimizes the error bounds in 2. For example, this could involve an additional nested training loop minimizing the error bound over a family of candidate kernels. Recent work in this area has shown promising results using kernel flows, a variant of cross-validation [16, 17].
@@ Appendix A RKHS Spaces, Optimal GP Recovery, and the Representer Theorem
$\begin{split}\ell(\mathbf{X};\mathbf{Y})&=\left\lVert f^{*}\right\rVert_{K}+\frac{1}{\sigma_{\epsilon}^{2}}\left\lVert f^{*}(\mathbf{X})-\mathbf{Y}\right\rVert_{2}^{2}\\ &=\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}+\frac{1}{\sigma_{\epsilon}^{2}}\left\lVert\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}-\mathbf{Y}\right\rVert_{2}^{2}\\ &=\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}\\ &\qquad\qquad+\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{K}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}\\ &\qquad\qquad-\frac{2}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}+\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{Y}\\ &\stackrel{{\scriptstyle(i)}}{{=}}\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}\\ &\qquad\qquad+\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}(\mathbf{K}+\sigma_{\epsilon}^{2}I-\sigma_{\epsilon}^{2}I)\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}\\ &\qquad\qquad-\frac{2}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}+\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{Y}\\ &=\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}\\ &\qquad\qquad+\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}-\mathbf{Y}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}\\ &\qquad\qquad-\frac{2}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}+\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{Y}\\ &=-\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{K}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}+\frac{1}{\sigma_{\epsilon}^{2}}\mathbf{Y}^{T}\mathbf{Y}\\ &\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{\sigma_{\epsilon}^{2}}\left(-\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I-\sigma_{\epsilon}^{2}I)(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}+\mathbf{Y}^{T}\mathbf{Y}\right)\\ &=\frac{1}{\sigma_{\epsilon}^{2}}\left(-\mathbf{Y}^{T}\mathbf{Y}+\sigma_{\epsilon}^{2}\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y}+\mathbf{Y}^{T}\mathbf{Y}\right)\\ &=\mathbf{Y}^{T}(\mathbf{K}+\sigma_{\epsilon}^{2}I)^{-1}\mathbf{Y},\\ \end{split}$ (70)
@@ Appendix B Block matrix formulation
for $i\in\mathcal{E}_{\text{un}}$, where $K_{i}$ is the kernel learned for edge $i$ [^7] Recall that we learn a lengthscale parameter for each edge, thus defining a unique kernel on each edge of the graph. and $\mathbf{u}_{i,j}$ denotes the $j$th datapoint for the endpoints of edge $i$.

### v2

@@ 5.2 Subsurface fracture network problem
Discrete fracture networks serve as preferential pathways for fluid flow and heat transport in subsurface systems. Understanding the dynamics of such systems is essential for applications such as carbon sequestration, geothermal energy, and hydraulic fracturing. However, observations of such systems are typically both sparse and noisy, rendering traditional approaches insufficient for proper study. Surrogate modeling with deep neural networks has been proposed as one potential approach to studying fracture network properties [song_surrogate_2024]. In the current work, we explore Dirichlet-to-Neumann maps on graphs as an alternative approach.
We use the synthetic 2-D fracture network dataset of [song_surrogate_2024], generated by a physics-based Darcy-flow solver. The fracture geometry is generated according to the procedure detailed in [delphine_fractures, song_surrogate_2024], and is characterized by the percolation parameter, $p=9$, and aperture, $a=1.5$. Fracture segments are designated as edges, and fracture intersections and terminations are designated as vertices. The resulting graph, depicted in Figure 3, has $V=107$ vertices and $E=130$ edges, of which 17 are designated boundary boundary entities (nine injection sites with high hydraulic head and eight extraction sites with low hydraulic head).
@@ 5.3 Arterial blood flow problem
The high computational cost of building high fidelity models with full field patient-specific data makes arterial blood flow another appealing test case for this approach. Arterial blood flow is a challenging test case because it exhibits both nonlinear dynamics and potentially complicated geometry. Our dataset consists of arterial graphs generated as training data for the GNN reduced order model developed in [pegolotti_learning_2023]. Data generation involves first running the 3D hemodynamic simulation SimVascular [updegrove2017simvascular], then converting the data to a 1D representation and ultimately into a graph. The full data generation process is described in Section 4 of [pegolotti_learning_2023].
@@
Because the approach described in [pegolotti_learning_2023] calculated the artery aperture using a cutting-plane method, the dataset can be nonphysical (and hence, non-conservative) in areas near intersections, since a single plane may cut more than one arterial branch. For this reason, when we consider results on the interior of the graph, we focus on the “branch” vertices where this data fidelity issue does not arise (highlighted in Figure 5 Panel B).
@@
The cutting-plane artifact in [pegolotti_learning_2023] breaks local mass conservation near junctions; our model, by construction, enforces a globally conservative flow field that minimizes the RKHS norm subject to the boundary data. Consequently, the largest discrepancies between the predictions and truth appear exactly in those non-conservative regions, as visualized by the blue and red shading in the rightmost panel of Figure 6. For branch vertices—where the true data are conservative—errors are small and are captured by the estimated uncertainty bound. This is because our approach constructs a conservative model obeying the boundary data regardless of whether the ground truth data is actually locally conservative everywhere. It is also important to note that we expect our estimated uncertainty bound to underestimate the error on the interior due to the fact that the bound assumes the GP input values on the interior are accurate, when they are actually inferred estimates (see Section 4.4).
@@ 6.1 Future work
We assumed small Gaussian noise in our observations. The effect of this assumption is the regularization of the kernel matrix with a small diagonal term $\sigma_{\epsilon}^{2}I$, referred to as a “nugget.” For simplicity, we have assumed a uniform value for $\sigma_{\epsilon}^{2},$ but in reality the spectral properties of the different edges of our graph (and thus the different blocks of the kernel matrix) may differ widely. An adaptive nugget term, such as that proposed by [chen_solving_2021], could more optimally balance accuracy and stability across the different edges of the graph.
Our current treatment does not account for the uncertainty stemming from the inferred input values to the GPs. We assume that the interior vertex values are unknown and estimated via the procedure described in Section 4.4, but the error estimates provided in Section 4.5 do not account for this. To incorporate this additional source of uncertainty, future work could explore linearizing the problem around the MAP, for example via a Laplace approximation as in [chen_gaussian_2024].
Finally, future work could explore learning the kernel $K(\cdot,\cdot)$ that minimizes the error bounds in 2. For example, this could involve an additional nested training loop minimizing the error bound over a family of candidate kernels. Recent work in this area has shown promising results using kernel flows, a variant of cross-validation [learningkernels2, learningkernels].
@@ Appendix A RKHS Spaces, Optimal GP Recovery, and the Representer Theorem
[…]
@@ Appendix B Block matrix formulation
for $i\in\mathcal{E}_{\text{un}}$, where $K_{i}$ is the kernel learned for edge ii [^7] Recall that we learn a lengthscale parameter for each edge, thus defining a unique kernel on each edge of the graph. and $\mathbf{u}_{i,j}$ denotes the $j$th datapoint for the endpoints of edge $i$.